# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

### v1

@@ 8.3. Ethical Concerns and Risk of Using LLM for Persuasion
Meanwhile, privacy is another critical concern since our system detects users’ location (as general categories) and captures their smartphone usage data. MindShift employs a commercial API, transmitting users’ data to a third party. Although we intentionally designed the prompt to avoid including any identifiable information, there is still the risk of revealing information about their behavior and mental states. One solution for future study is to leverage open-sourced LLMs (such as LLaMA2 (Touvron et al., 2023) or PaLM2 (Anil et al., 2023)) so that user’s data can be appropriately handled and encrypted by ourselves instead of a third party.
@@ 8.4. Limitations
There are a few limitations in our study. First, our experimental user group is limited to college students, which limits the applicability of the results. Future studies can expand the sample size and involve more diverse user groups. Second, our field experiment is short. A five-week deployment cannot reveal the longitudinal effect of such an intervention technique. Moreover, if our time and monetary budget allow, our experiment design can be improved by making the Baseline another two-week intervention session for a more fair comparison. Third, our current method for detecting habitual use and mental states relies on self-reporting, which increases users’ burden. We only consider initial habitual use and do not consider the shifts in user purposes during app usage. As we mentioned in Section 8.1, future work can explore automatic intent and mental state detection, and the data collected in this study can serve as a starting point for machine learning training models. Finally, our study utilizes GPT-3.5 as a large language model, and its performance is still unstable. Future work can explore more lighted weighted and robust LLMs for local deployment, which can address the concerns mentioned in Section 8.3 to some extent.
@@ 9. CONCLUSION
In this paper, we propose MindShift, a mental-based persuasion intervention technique powered by LLM that aims to reduce problematic smartphone use. We conduct a Wizard-of-Oz study and an interview study to explore the mental states behind problematic smartphone use: stress, boredom, inertia, and design four persuasion strategies: understanding, comforting, evoking, and scaffolding habits. MindShift (1) collects users’ purpose, usage behavior, physical context, mental states, goals&habits, (2) uses the persuasion strategy we design, (3) leverages LLMs to generate dynamic, personalized persuasion messages. Through a five-week within-subjects user experiment (N=25), we compare three intervention techniques (MindShift, MindShift-Simple, Baseline). Our results show that MindShift significantly improves the acceptance rate of generated persuasion content (8.1%) compared with MindShift-Simple. Two versions of MindShift decrease overall app opening frequency (12.1-14.4%) and usage duration (2.4-9.8%) and are particularly effective in reducing habitual use. Moreover, two versions of MindShift significantly reduced SAS scores (25.8-34.7%) and increased self-efficacy scores (10.4-10.7%). Finally, users’ subjective comments also confirm a perceived reduction in smartphone dependency and a willingness to continue to use MindShift. Our work provides valuable insights into the mental states behind problematic smartphone use and the effectiveness of LLMs-powered persuasion for smartphone intervention.

### v2

@@ 8.3.3. Ethical Concerns and Risk of Using LLMs for Behavior Change
Moreover, privacy is another critical concern since the detection of users’ physical and psychological context data is needed. MindShift employs a commercial API from OpenAI, transmitting users’ data to a third party. Although we intentionally designed the prompt to avoid including any identifiable information, there is still the risk of revealing information about their behavior and mental states. One solution for future study is to leverage open-sourced LLMs (such as LLaMA2 (Touvron et al., 2023) or PaLM2 (Anil et al., 2023)) so that user’s data can be appropriately handled and encrypted by ourselves instead of a third party.
@@ 8.4. Limitations
Our study has a few limitations. First, our experimental user group is limited to young adults, limiting result applicability. Future studies can expand the sample size and involve more diverse user groups. Second, our field experiment is short. A five-week deployment cannot reveal the longitudinal effect of such an intervention technique. Moreover, if our time and monetary budget allow, our experiment design can be improved by making the Baseline another two-week intervention session for a more fair comparison. Third, the validity and reliability test of mental state report questions needs further improvement. Testing convergent and discriminant validity, recruiting more samples for reliability tests, and using statistical methods to evaluate consistency are areas for enhancement. Additionally, our current method for detecting habitual use and mental states relies on self-reporting, increasing users’ burden. We only consider initial habitual use upon users unlocking phones, neglecting shifts in user purposes during app usage. As we mentioned in Sec. 8.3.2, future work can explore automatic intent and mental state detection, and the data collected in this study can serve as a starting point for machine learning training models. Finally, our study utilizes GPT-3.5 as a large language model, and its performance is still unstable. Future work can explore more lighted weighted and robust LLMs for local deployment, which can address the concerns mentioned in Sec. 8.3.3 to some extent.
@@ 9. CONCLUSION
This paper introduces MindShift, a mental-based persuasion intervention technique powered by LLM designed to mitigate problematic smartphone use. We conducted a Wizard-of-Oz study and an interview study to explore the mental states behind problematic smartphone use: stress, boredom, inertia, and designed four persuasion strategies: understanding, comforting, evoking, and scaffolding habits. MindShift (1) collects users’ usage intent, usage behavior, physical context, mental states, goals&habits, (2) uses the persuasion strategies we design, (3) leverages LLMs to generate dynamic, personalized persuasion messages. Through a five-week within-subjects user experiment (N=25), we compared three intervention techniques (MindShift, MindShift-Simple, Baseline). MindShift outperforms MindShift-Simple and Baseline, improving acceptance rates (4.7-22.5%) and reducing app usage (7.4-9.8%). Notably, MindShift and MindShift-Simple significantly reduce SAS scores (34.7-25.8%) and increase self-efficacy scores (10.7-10.4%). Finally, users’ subjective comments also confirm a perceived reduction in smartphone dependency and a willingness to continue to use MindShift. Our work provides valuable insights into the mental states behind problematic smartphone use and the effectiveness of LLMs-powered persuasion for smartphone intervention.
Appendix
Table 4. Persuasive Messages in WoZ Study. Four types of persuasive messages delivered in WoZ study and their examples.
Types Examples
Usage Notice ”You have used Wechat for 2 hours and 25 minutes today. Put down your phone please!”
Practical Guidance ”Are you still using WeChat? Have you completed the task of analyzing data today?”
Encouragement ”You have only spent 2 hours on your phone today, that’s excellent! Keep up the good $\sim$
Deterrent ”Using the phone before bedtime can affect the quality of your sleep.”
Table 5. Takeaways from WoZ & semi-structured interview studies. (E) shows messages sent by experimenters, (W) demonstrates participants’ reflection quotes in the WoZ study, and (S) means participants’ quotes in the semi-structured interview study.
WoZ study
Types of smartphone use Representative quote(s)
Instrumental use
(not to be intervened)
”You’ve already spent one and a half an hour on WeChat today. Please put your phone down and focus on other aspects of life. (E)”
”I felt a bit resentful because I was using WeChat to manage my affairs, rather than idly wasting time. (W1)”
Instrumental use - relaxation
(not to be intervened)
”Please stop browsing Zhihu and engage in more meaningful activities. (E)”
”I don’t agree. Finding joy in browsing Zhihu constitutes meaning for me. (W5)”
Habitual use
(to be intervened)
”You have used Zhihu for 2 hours today. Think about what else you have to do tonight. (E)”
”Thanks for this suggestion. I always failed to control myself to open Zhihu. (W10)”
Factors affecting persuasion effectiveness Representative quote(s)
Mental states
”When you find yourself with idle time, consider engaging in meaningful activities such as reading, writing, or drawing. (E)”
”It correctly identified my state of not knowing what to do and offered sensible advice. (W11)”
Personal goals
”Your WeChat session has lasted 10 minutes. Please set aside your device to alleviate eye strain. (E)”
”Keeping the eyes healthy is one thing I really care about, so I like this advice. (W12)”
Contextual information
”The afternoon is a good time for studying. Don’t spend too much time on your phone. (E)”
”Afternoon is indeed my study time during which I should improve my efficiency. It is right. (W4)”
Semi-structured interview study
Mental states related to habitual use Representative quote(s)
Boredom
”I find myself instinctively reaching for my phone in search of mental stimulation when doing simple assignments light on cognitive engagement.” (S1)
Stress
”One day, work wasn’t progressing well and I was so frustrated that I unlocked my phone for a quick view to ease my mood.” (S9)
Inertia
”Upon returning home after work, I intended to transition back into a focused state for reading or other activities but struggled to shift from a relaxed state. At that point, my phone was the tool for procrastination.” (S2)
Activity engagement states Representative quote(s)
Engaging in activities
”I was reluctant to start handling this challenging work that I scrolled my phone screen anxiously.” (S2)
Not engaging in activities
”After getting off work and returning home, I collapse on the sofa and binge-watch Tiktok for one to two hours.” (S9)
Figure 12. Interaction Process of Three Intervention Techniques. The two red process blocks illustrate the differences between the three intervention techniques.
This flowchart outlines the interaction process of three intervention techniques. Red process blocks illustrate the differences between the three intervention techniques. It starts with the user opening a restricted app and reporting their intent. Then comes to a red block to determine whether it is baseline. If the intervention technique is not the baseline, the user’s habitual intent is assessed. If the intent is habitual, a persuasion intervention is triggered. Then comes to a red block to determine whether it is MindShift. Depending on whether the technique is MindShift or a simpler version, the strategy and context or just the context are included. The intervention continues based on acceptance: if accepted, the user is prompted to quit the app; if not, the process checks if the accumulated usage time is over 2 minutes or if the current round has reached the number of strategies planned. If neither condition is met, the intervention ends.
Figure 12. Interaction Process of Three Intervention Techniques. The two red process blocks illustrate the differences between the three intervention techniques.